# The Curious Case of In-Training Compression of State Space Models

### v3

@@ F.1.1 Differentiable Hankel Nuclear Norm Regularization
For a diagonal state-transition matrix ${\bm{A}}_{\bm{\theta}}=\mathrm{diag}({\bm{\lambda}}_{\bm{\theta}})$, the Gramians admit closed-form solutions. The discrete-time controllability and observability Lyapunov equations,
${\bm{A}}_{\bm{\theta}}{\bm{P}}_{\bm{\theta}}{\bm{A}}_{\bm{\theta}}^{T}-{\bm{P}}_{\bm{\theta}}+{\bm{B}}_{\bm{\theta}}{\bm{B}}_{\bm{\theta}}^{T}$ $=0,$ (23)
${\bm{A}}_{\bm{\theta}}^{T}{\bm{Q}}_{\bm{\theta}}{\bm{A}}_{\bm{\theta}}-{\bm{Q}}_{\bm{\theta}}+{\bm{C}}_{\bm{\theta}}^{T}{\bm{C}}_{\bm{\theta}}$ $=0,$ (24)
@@
$({\bm{P}}_{\bm{\theta}})_{ij}$ $=\frac{({\bm{B}}_{\bm{\theta}}{\bm{B}}_{\bm{\theta}}^{T})_{ij}}{1-\lambda_{{\bm{\theta}},i}\lambda_{{\bm{\theta}},j}},$ (25)
$({\bm{Q}}_{\bm{\theta}})_{ij}$ $=\frac{({\bm{C}}_{\bm{\theta}}^{T}{\bm{C}}_{\bm{\theta}})_{ij}}{1-\lambda_{{\bm{\theta}},i}\lambda_{{\bm{\theta}},j}}.$ (26)
@@ F.2 Knowledge Distillation
The effective knowledge distillation loss for classification $\mathcal{L}$ is a superposition of the cross-entropy loss between the student and the targets and the Kulback-Leibler divergence between the teacher and the student with temperature scaling:
$\mathcal{L}=(1-\alpha)\mathcal{H}(y,\sigma(z_{s}))+\alpha T^{2}D_{KL}(\sigma(z_{t}/T)||\sigma(z_{s}/T))$

### v4

@@ F.1.1 Differentiable Hankel Nuclear Norm Regularization
For a diagonal state-transition matrix ${\bm{A}}_{\bm{\theta}}=\text{diag}({\bm{\lambda}}_{\bm{\theta}})$, the Gramians admit closed-form solutions. The discrete-time controllability and observability Lyapunov equations,
${\bm{A}}_{\bm{\theta}}{\bm{P}}_{\bm{\theta}}{\bm{A}}_{\bm{\theta}}^{\top}-{\bm{P}}_{\bm{\theta}}+{\bm{B}}_{\bm{\theta}}{\bm{B}}_{\bm{\theta}}^{\top}$ $=0,$ (23)
${\bm{A}}_{\bm{\theta}}^{\top}{\bm{Q}}_{\bm{\theta}}{\bm{A}}_{\bm{\theta}}-{\bm{Q}}_{\bm{\theta}}+{\bm{C}}_{\bm{\theta}}^{\top}{\bm{C}}_{\bm{\theta}}$ $=0,$ (24)
@@
$({\bm{P}}_{\bm{\theta}})_{ij}$ $=\frac{({\bm{B}}_{\bm{\theta}}{\bm{B}}_{\bm{\theta}}^{\top})_{ij}}{1-\lambda_{{\bm{\theta}},i}\lambda_{{\bm{\theta}},j}},$ (25)
$({\bm{Q}}_{\bm{\theta}})_{ij}$ $=\frac{({\bm{C}}_{\bm{\theta}}^{\top}{\bm{C}}_{\bm{\theta}})_{ij}}{1-\lambda_{{\bm{\theta}},i}\lambda_{{\bm{\theta}},j}}.$ (26)
@@ F.2 Knowledge Distillation
The effective knowledge distillation loss for classification $\mathcal{L}$ is a superposition of the cross-entropy loss between the student and the targets and the Kullback-Leibler divergence between the teacher and the student with temperature scaling:
$\mathcal{L}=(1-\alpha)\mathcal{H}(y,\sigma(z_{s}))+\alpha T^{2}D_{\mathrm{KL}}(\sigma(z_{t}/T)\|\sigma(z_{s}/T))\,.$